INAF LBC+MODS queue observing, 2013 DATA SUMMARY January

Observer: M. Faccini, R. Speziali, F. Cusano
Telescope Operator: D. Gonzales-Huerta, G. Bechetti
Support Astronomer: M. Edwards, O. Kuhn
Instrument: LBC+MODS

January 10-11

Dome closed due to Clouds and Wind over 26m/s (weather forecasted snow 90%)

Project Objects Filter ExpTime Notes
         

January 11-12

Just few clouds at west horizon at sunset, sky is clear. Dewpoint at -16\xB0C and external temperature at - 15\xB0C
There is a difference about 12/13 degres between primary mirrors and external temperature
We have to wait changes in weather conditions.
After few hours the external temperature falls down to -19\xB0C, the mirrors temperature is still far from the external temperature. At 3:30 the TO decides to not open the dome for this night ( we are almost at -20\xB0C)

January 12-13

Dome CLOSED due to low temperature, almost -20\xB0C and difference between external and mirrors temperature, about 14\xB0C

January 13-14

Heavy clouds on some part of the sky at sunset. We open dome at 12:40UT.
External temperature is about -16 degres, difference between primary morrors temperature and external temperature is about 8 degres
Observation started with seeing 1.8"
We close the Dome due to technical problem and low temperature -17, the dome was stucked and we close the dome by "hands"
Project_RUN Objects Filter ExpTime Notes
13 A3-A4 Battaglia LGS3 V and I 30'  
27D Giallongo CLJ0956+41 g and r 1h some images elongated and with bad seeing

January 14-15

At sunset temperature is about -17/-18, clouds and few snow.
Weather seems variable as foreseen.
Last night problems with MODS still not solved, at 0:30 UT technicians are still working on LBC-red to solve the problem about winter position of the cameras.
at the beginning of the night the sky is clear and ext temperature is about -19/-20 at less than 1 degree from dewpoint temperature. We can not open the dome.
Around 7 UT temperature decrease to -17 and dewpoint is at -25, but we have snow on the roof. The Dome will remain close due to snow on the roof.

January 15-16

At sunset temperature is about -10, some cirrus far away, wind gust at ~ 15 to 23 m/s
Last nights problems with MODS still not solved
After opening the dome (03:15 UT) we have problems with focus, there is an intra focus on red channel. impossible to solve sending manual offset. Crew have to move the red camera manually.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Giallondo 27D CLJ0956+41 g-sloan, r-sloan 1.5h Total integration time 3h,but 1.5h in the spec for the seeing
Sky flat   g-sloan, r-sloan    
Standard SA98 Us, Y, g, r    

January 16-17

We had problems to focus the telescope at the beginning of the night (almost one hour lost), at the beginning of the night outside temperature 3-4 degree C while mirror temperature -4. Average seeing of the night 1 arcsec. We carried out the program Giallondo 27D and part of Grazian 30A.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Giallondo 27D CLJ0956+41 g-sloan,r-sloan 2h total exposure time >2h, but some images were out of focus
Grazian 30A GOODSN STD-Uspec, r-sloan 2h total exposure time >2h, but some images were out of focus
Standard SA98 STD-Uspec,Y,g,r    
Sky flats   V-Johnson, I-Johnson    

January 19-20

Clear sky, seeing 0.8 arcsec, moon at the first quarter. We started with LBC. Problem at the beginning of the night, stars elongated in the blue channel, we tried to fix it manually (more than one hour), but it did not work. We switch to MODS.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Annibali 23B hd74377 Dual grating 2x1s Lick standard COMPLETED
Annibali 23B hd84937 Dual grating 1s Lick standard COMPLETED
spectrophotometric standard GD71 Dual grating 3x180s  
spectrophotometric standard Hiltner 600 Dual grating 3x60s  
spectrophotometric standard G191-B2B Dual grating 3x90s  
Annibali 23B NGC4449 Mos dual grating 2h science target
Fritz 4A LensGG243 Mos dual grating 2h science target COMPLETED
Annibali 23B Feige66 Dual grating 1x120s  
Gilli 18D Feige66 Red grating 3x90s  
Gilli 18D cxoj12 Red grating 4x900s science target;file 0058 wrong filter; 0059 correct filter but header wrong
Annibali 23B hd160693 Dual grating 1s Lick standard

January 20-21

At the beginning of the night thick clouds at the horizon, cirrus all over in the sky, moon 63% illumination, seeing 0.8. Starting with LBC. Problems with LBC, elongated stars in the blue channel. Trying to solve the problem. Mostly clear in the middle of the night. Switch to MODS.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Annibali 23B NGC 4449 dual grating 2h science target COMPLETED
Annibali 23B HD108177 dual grating 1s Lick standard COMPLETED
Annibali 23B Feige 66 dual grating   spec standard
Glli 18D Feige 66 red grating   spec standard
Glli 18D cxoj12 red grating 15m science target COMPLETED
Annibali 23A NGC 4449 Pn dual grating 2h science target

-- MarcoFaccini - 11 Jan 2013
Topic revision: r6 - 29 Jan 2013, RobertoSpeziali
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback